# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 6.3.2. DP-based ILFPPM
The privacy norm of DP is also utilized to protect privacy in indoor localization. The proposed methods lie on one of the Centralized Differential Privacy (CDP), Local Differential Privacy (LDP), or Geo-indistinguishability. CDP is a framework for protecting the privacy of individual data points in a centralized dataset. In this approach, a trusted server applies DP mechanisms to the entire dataset to ensure that the analysis or queries performed on the data do not reveal sensitive information about any individual data point. In LDP, on the other hand, individual data contributors apply privacy mechanisms to their data before sending it to a central server or aggregator. Each data point is perturbed or obfuscated in such a way that, when aggregated with other data points, the perturbed data still provides statistically valid information for analysis, but does not compromise the privacy of the individuals. This difference between CDP and LDP is plotted in Fig. 10. In addition, geo-indistinguishability is a DP-based technique used to protect the privacy of individuals’ locations by adding noise to local data, thus leading to obfuscating the real location.
Table 3 provides a list of ILFPPM based on DP. To illustrate, authors in (Li et al., 2016) utilize DP to inject noise into participants’ fingerprint data to protect location privacy. In contrast, the study conducted by (Zhu et al., 2017) extends DP and proposes a data obfuscation mechanism to generalize and specialize users’ fingerprints. Another study, as presented in (Sadhu et al., 2017), designs a localization platform relying on Wi-Fi RSS values and introduces data perturbation and randomization techniques grounded in DP. The aforementioned works concentrate on privacy preservation within specific types of indoor localization. The work described in (Zhao et al., 2018) introduces a privacy-preserving paradigm-driven framework for indoor localization utilizing local differential privacy. This approach is motivated by the observation that many IPS adhere to a common two-stage localization paradigm. Moreover, the study presented in (Kim and Jang, 2019) combines local differential privacy and optimal data encoding to disturb users’ data. Similarly, authors in (Wang et al., 2018) propose a DP-based privacy-preserving indoor localization scheme that introduces noise to the data, such as RSSI, CSI, and so forth, which increases the likelihood of significant localization errors and lowers the standard of localization services. The work in (Zhang et al., 2022) applies fusion of signal gathered from multiple wireless technologies (e.g., WiFi and BLE) with differentially private fingerprint fusion semi-supervised extreme learning machine for indoor localization in the edge computing, called Adp-FSELM. Another study introduces a novel privacy-aware framework for aggregating indoor location data employing the LDP technique. In this approach, the user location data is locally transformed on the user’s device and subsequently sent to the aggregator (Navidan et al., 2022).
@@
Employing other DP-based solutions, the authors in (Min et al., 2022) initially introduce 3D Geo-indistinguishability for the application of indoor localization. This is followed by the addition of reinforcement learning for the sake of semantic location privacy in (Min et al., 2023). These papers assume that the user has access to her location and sends her location for obtaining services. From another perspective, (Fathalizadeh et al., 2023) introduces two methods for applying geo-indistinguishability based on RSS vectors without having access to the location.
Figure 11. An overview of FL in indoor location fingerprinting systems
6.4. Federated Learning
@@
Here, we briefly explain the FL mechanism and then discuss the FL-based ILFPPM.
6.4.1. Federated Learning mechanism
@@
Google debuted FL, a privacy-preserving distributed ML approach, in 2016 (McMahan et al., 2017). This method trains an algorithm on several decentralized edge devices or local servers without requiring them to exchange raw data, in contrast to traditional centralized machine learning techniques that load all local datasets onto a single server. Initially, a machine learning model is trained by a centralized server and this model is then sent to local devices that hold their own private data. On these respective devices, the local models are trained using their private data, as shown in Fig. 11. The training process typically involves using local computational resources and updates the model based on the local data. After training, the locally updated models send their changes (e.g., parameters or gradients) back to the centralized server, which aggregates these updates. Finally, the server updates the global model using an aggregation method such as FedAvg and FedAmp (Hsu et al., 2019; Wu et al., 2021a). FedAvg is one of the most common aggregation methods for FL, where the central server combines the gradients sent by the users using the below equation and updates the global model.
(4) $w_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}w^{k}_{t},$
@@
Table 4. Studies on FL-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Attack Model (Liu et al., 2019) 2019 DNN AutoEncoder Private Coms. Overhead Loc. Error Loc. Priv. Attack (Li et al., 2020) 2020 SAEC FedAvg UJIIndoorLoc Loc. Error Loc. Priv. Attack (Ciftler et al., 2020) 2020 MLP FedAvg UJIIndoorLoc Loc. Error - (Wu et al., 2021b) 2021 MLP FedAvg FedAmp FedAmp-fusion UJIIndoorLoc Loc. Error - (Wu et al., 2021c) 2021 MLP FedAvg DP UJIIndoorLoc Guangzhou Xinguang shopping mall Loc. Error Noise Effect Inference Attack Loc. Priv. Attack (Tasbaz et al., 2022) 2022 kNN, MLP Random Forest FedAvg UJIIndoorLoc CRI Accuracy Run Time Loc. Priv. Attack Probability-based Attack (Etiabi et al., 2022) 2022 DNN Knowledge distillation Federated distillation Simulated UJIIndoorLoc Coms. Overhead Loc. Error - (Cheng et al., 2022) 2022 MLP FedAvg UJIIndoorLoc Accuracy Loc. Error - (Wu et al., 2022b) 2022 MLP FedAvg Knowledge distillation Mixture of Experts UJIIndoorLoc Guangzhou Xinguang shopping mall Loc. Error Inference Attack Loc. Priv. Attack (Nagia et al., 2022) 2022 CNN, LSTM FedProx Private Loc. Error - (Park et al., 2022) 2022 NN Monte Carlo Dropout UJIIndoorLoc Loc. Error Time Cost - (Wu et al., 2022a) 2022 GNN Federated Graph Learning Guangzhou Xinguang shopping mall Loc. Error - (Gao et al., 2023) 2023 CNN, DNN AutoEncoder UJIIndoorLoc Accuracy Loc. Error Coms. Overhead Inference Attack Loc. Priv. Attack (Kumar et al., 2023) 2023 CNN, LSTM - Private Accuracy - (Gufran and Pasricha, 2023) 2023 MLP AutoEncoder FedHIL Private Coms. Overhead Loc. Error Inference Attack (Guo et al., 2022) 2023 CNN Cloud Feature Extractor Homomorphic Encryption Simulated Coms. Overhead Loc. Error Inference Attack Probability-based Attack (Varma et al., 2023) 2023 kNN Discrete Coordinate Encryption Federated KNN Private Loc. Error Energy Consumption - (Yan et al., 2024) 2024 CNN Inner Production private Accuracy Loc. Error Data Priv. Attack
6.4.2. FL-based ILFPPM
@@
As FL enables training a centralized model across decentralized devices or servers holding local data samples, without exchanging the data samples, it has been applied to IPS. IPS often involves collecting location data from various devices and using FL allows these devices to collaboratively learn a positioning model without sharing the raw location data. Accordingly, instead of sending all raw data to the central server, only model updates or aggregated information need to be communicated, thus reducing the amount of data transmitted over the network. Since indoor locations vary very widely, this approach is very useful when updating the model. So under the FL setting, indoor positioning models can be trained using data from various users in different environments without having to centralize that data. This helps protect the privacy of users while still allowing for model improvements and leads to a more accurate and robust IPS. Despite the benefits of applying FL in IPS, it also poses technical challenges including ensuring model convergence despite the heterogeneity of users’ data and addressing bias introduced by variations in local data. Table 4 reports ILFPPM studies in the FL setting. Most studies focus on the RSS data, but CSI is also taken into consideration in some papers. The proposed aggregation methods mostly are FedAvg and FedAmp, followed by widely used MLP and CNN learning algorithms.
6.5. Overall Comparison

### v2

@@ 5.5.2 Comparative analysis of DP-based techniques
The privacy norm of DP is also utilized to protect privacy in indoor localization. The proposed methods lie on one of the Centralized Differential Privacy (CDP), Local Differential Privacy (LDP), or Geo-indistinguishability. CDP is a framework for protecting the privacy of individual data points in a centralized dataset. In this approach, a trusted server applies DP mechanisms to the entire dataset to ensure that the analysis or queries performed on the data do not reveal sensitive information about any individual data point. In LDP, on the other hand, individual data contributors apply privacy mechanisms to their data before sending it to a central server or aggregator. Each data point is perturbed or obfuscated in such a way that, when aggregated with other data points, the perturbed data still provides statistically valid information for analysis, but does not compromise the privacy of the individuals. This difference between CDP and LDP is shown in Fig. 10. In addition, geo-indistinguishability is a DP-based technique used to protect the privacy of individuals’ locations by adding noise to local data, thus leading to obfuscating the real location.
Table 5 provides a list of ILFPPM based on DP. One of the earliest implementations of DP-based ILFPPM is presented by [95], which integrates CDP with Paillier encryption for protecting kNN-based localization. This method focuses on communication overhead and location error as performance metrics, showing its potential in a semi-trusted adversary model. However, location privacy attacks remain a critical vulnerability, indicating the need for more refined adversary models to address advanced threats. In contrast, [61] and [96] employed LDP mechanisms, offering a more decentralized privacy guarantee. [61]’s work applies two-step classification for building-level indoor localization using real-world datasets such as Rutgers University. [96], on the other hand, adopts Delaunay triangulation and range generalization, focusing heavily on reducing location error. Although both approaches show promise, they lack explicit adversary models, which makes it difficult to gauge their robustness against more sophisticated privacy attacks.
The aforementioned works concentrate on privacy preservation within specific types of indoor localization. The work described in [97] introduces a privacy-preserving paradigm-driven framework for indoor localization utilizing LDP. This approach is motivated by the observation that many IPS adhere to a common two-stage localization paradigm. As LDP evolved, [60] introduced randomized response as an LDP mechanism, showing improvements in reducing error rates under fully malicious adversaries. This study further demonstrated the method’s efficacy in countering location privacy and probability-based attacks using Sangmyung University data. However, despite its strengths, randomized response techniques often introduce noise, potentially increasing the communication overhead or compromising data utility. Moreover, the study presented in [98] combines local differential privacy and optimal data encoding to disturb users’ data. Similarly, authors in [57] propose a DP-based privacy-preserving indoor localization scheme that introduces noise to the data, such as RSSI, CSI, and so forth, which increases the likelihood of significant localization errors and lowers the standard of localization services.
The work in [100] applies fusion of signal gathered from multiple wireless technologies (e.g., WiFi and BLE) with differentially private fingerprint fusion semi-supervised extreme learning machine for indoor localization in the edge computing, called Adp-FSELM. It incorporates edge-cloud collaboration to optimize communication overhead and run time, while maintaining a balance between location error and privacy protection in the presence of fully malicious adversaries. Another study in  [99] introduces a novel privacy-aware framework for aggregating indoor location data employing the LDP technique. It combines multiple privacy-preserving techniques like local hashing, unary encoding, and RAPPOR, using datasets like JUIndoorLoc and CRI. In this work, the user location data is locally transformed on the user’s device and subsequently sent to the aggregator. This comprehensive approach is particularly effective against location and data privacy attacks, showcasing the versatility of LDP in practical scenarios.
@@
Employing other DP-based solutions, the authors in [101] initially introduce 3D Geo-indistinguishability for the application of indoor localization. This is followed by the addition of reinforcement learning for the sake of semantic location privacy in [103]. These papers assume that the user has access to her location and sends her location for obtaining services. From another perspective, [102] introduces two methods for applying geo-indistinguishability based on RSS vectors without having access to the location. These studies strongly emphasize balancing location privacy and QoS loss while addressing increasingly complex attack models like semantic, inference, and context-linking attacks. The trade-off between privacy guarantees and QoS is a recurring challenge, indicating the need for further refinement of DP-based ILFPPM to meet real-world demands without sacrificing data utility.
5.5.3 Discussion on employing DP-based techniques
To implement DP-based techniques in indoor localization systems, the following guidelines will help ensure both privacy protection and operational efficiency:
Selection of DP model: When choosing between CDP and LDP, trust assumptions play a key role. CDP is suitable when a trusted centralized authority is available, as it adds noise to aggregated data at the server level, allowing for global optimization. However, in cases where users don’t fully trust the data aggregator or want to minimize reliance on a central entity, LDP is more appropriate. LDP ensures privacy at the user level by adding noise before data leaves the device, making it ideal for decentralized environments or scenarios where user control and anonymity are crucial. The choice depends on balancing privacy, trust, and system architecture.
Mitigating attack models: In environments vulnerable to fully malicious adversaries, such as those discussed by LDP and CDP, more advanced privacy mechanisms like randomized response, edge-cloud collaboration, or GeoInd are often necessary. These methods provide stronger resilience against a broader spectrum of attacks, including location privacy breaches, probabilistic attacks, and sophisticated inference-based threats. Randomized response allows individual users to inject randomness into their data, supporting privacy at the source level. GeoInd [102], by introducing geographic obfuscation, further shields user location data, making it harder for adversaries to infer sensitive information.
Balancing privacy and utility: Implementations must achieve a careful balance between privacy protection and data utility. LDP techniques commonly inject noise to enhance privacy, which can result in a reduction in data utility, often evaluated using metrics such as location error or QoS loss. It is crucial to select mechanisms that deliver the required level of privacy while ensuring that the noise introduced remains within acceptable limits for the localization accuracy. This balance allows for maintaining effective data usability and performance, enabling systems to function optimally without compromising user privacy. Additionally, evaluating various LDP methods can help identify those that minimize adverse impacts on utility, facilitating better overall system performance and user satisfaction.
Optimization of communication overhead: When designing real-time systems or large-scale deployments, it is vital to assess the communication overhead associated with the selected DP mechanism. Techniques like randomized encoding [98] and Laplace mechanisms [102] can introduce significant overhead, potentially impacting overall system performance and responsiveness. This added complexity may lead to slower data processing and increased latency, which are critical in applications requiring immediate feedback or high throughput. Therefore, carefully evaluating the trade-offs between privacy protection and operational efficiency is essential, especially in resource-constrained environments where CPU power, bandwidth, and energy are limited. By considering these trade-offs, developers can select a DP approach that safeguards user privacy while maintaining system functionality and responsiveness, ensuring optimal performance across various operational contexts.
5.6 Federated Learning in ILFPPM
@@
Here, we briefly explain the FL mechanism and then discuss the existing FL-based ILFPPM. Finally, we provide guidelines on employing FL in ILF systems.
5.6.1 Federated learning mechanism
@@
Google introduced FL, a privacy-preserving distributed ML approach, in 2016 [104]. This method trains an algorithm on several decentralized edge devices or local servers without requiring them to exchange raw data, in contrast to traditional centralized machine learning techniques that load all local datasets onto a single server. Initially, a machine learning model is trained by a centralized server and this model is then sent to local devices that hold their own private data. On these respective devices, the local models are trained using their private data, as shown in Fig. 11. The training process typically involves using local computational resources and updates the model based on the local data. After training, the locally updated models send their changes (e.g., parameters or gradients) back to the centralized server, which aggregates these updates. Finally, the server updates the global model using an aggregation method such as FedAvg and FedAmp [105],[106]. FedAvg is one of the most common aggregation methods for FL, where the central server combines the gradients sent by the users using the below equation and updates the global model.
$w_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}w^{k}_{t},$ (4)
@@
Figure 11: An overview of FL in ILF systems
Table 6: Studies on FL-based ILFPPM.
Paper Year Localization Privacy / Aggregation Method Dataset Metric Attack Model
[107] 2019 DNN AutoEncoder Private
Coms. Overhead
Loc. Error
Loc. Priv. Attack
[108] 2020 SAEC FedAvg UJIIndoorLoc Loc. Error Loc. Priv. Attack
[109] 2020 MLP FedAvg UJIIndoorLoc Loc. Error -
[110] 2021 MLP
FedAvg
FedAmp
FedAmp-fusion
UJIIndoorLoc Loc. Error -
[111] 2021 MLP
FedAvg
DP
UJIIndoorLoc
Guangzhou Xinguang
shopping mall
Loc. Error
Noise Effect
Inference Attack
Loc. Priv. Attack
[112] 2022
kNN, MLP
Random Forest
FedAvg
UJIIndoorLoc
CRI
Accuracy
Run Time
Loc. Priv. Attack
Probability-based Attack
[113] 2022 MLP FedAvg UJIIndoorLoc
Accuracy
Loc. Error
[114] 2022 MLP
FedAvg
Knowledge distillation
Mixture of Experts
UJIIndoorLoc
Guangzhou Xinguang
shopping mall
Loc. Error
Inference Attack
Loc. Priv. Attack
[115] 2022 CNN, LSTM FedProx Private Loc. Error -
[116] 2022 NN Monte Carlo Dropout UJIIndoorLoc
Loc. Error
Time Cost
[117] 2022 GNN Federated Graph Learning
Guangzhou Xinguang
shopping mall
Loc. Error -
[118] 2023 CNN, DNN AutoEncoder UJIIndoorLoc
Accuracy
Loc. Error
Coms. Overhead
Inference Attack
Loc. Priv. Attack
[119] 2023 CNN, LSTM - Private Accuracy -
[120] 2023 MLP
AutoEncoder
FedHIL
Private
Coms. Overhead
Loc. Error
Inference Attack
[121] 2023 CNN
Cloud Feature Extractor
Homomorphic Encryption
Simulated
Coms. Overhead
Loc. Error
Inference Attack
Probability-based Attack
[122] 2023 kNN
Discrete Coordinate Encryption
Federated KNN
Private
Loc. Error
Energy Consumption
[123] 2024 CNN Inner Production Private
Accuracy
Loc. Error
Data Priv. Attack
[124] 2024 DNN
Knowledge distillation
Federated distillation
Simulated
UJIIndoorLoc
Coms. Overhead
Loc. Error
[125] 2024 CNN FedAvg Private
Accuracy
Loc. Error
[126] 2024 MLP, LSTM
FedAvg
FedSGD
SPAWC
Coms. Overhead
Loc. Error
Inference Attack
[127] 2024
kNN
SVM
Transfer Learning
Meta-learning
UJIIndoorLoc
Tampere Uni
Uni of Minho
Loc. Error
Adaptation speed
5.6.2 Comparative analysis of FL-Based techniques
@@
As FL enables training a centralized model across decentralized devices or servers holding local data samples, without exchanging the data samples, it has been applied to IPS. IPS often involves collecting location data from various devices and using FL allows these devices to collaboratively learn a positioning model without sharing the raw location data. Accordingly, instead of sending all raw data to the central server, only model updates or aggregated information need to be communicated, thus reducing the amount of data transmitted over the network. Since indoor locations vary very widely, this approach is very useful when updating the model. So under the FL setting, indoor positioning models can be trained using data from various users in different environments without having to centralize that data. This helps protect the privacy of users while still allowing for model improvements and leads to a more accurate and robust IPS.
Table 6 reports ILFPPM studies in the FL setting. Most studies focus on the RSS data, but CSI is also taken into consideration in some papers [121]. The reviewed studies employ AutoEncoders [107], [118], [120] in order to have the dimensionality reduction and privacy preserving data sharing when users can share encoded data. The most used aggregation method is also Federated Averaging (FedAvg) [108], [109], [114]. More advanced approaches, such as Federated Proximal (FedProx) [115], Federated Graph Learning [117], and Federated Meta-Learning [127], have been explored for specialized tasks. Additional methods, such as Monte Carlo Dropout [116] and Federated Knowledge Distillation [124], [114], represent further advancements within the FL paradigm. These methods emphasize the increasing diversity of privacy-enhancing techniques specified to different attack models.
In terms of datasets and performance metrics, the UJIIndoorLoc dataset is the most frequently used for evaluating these techniques, providing a consistent basis for measuring localization error and communication overhead [108], [112], [118]. Private datasets [107], [115], [121] have also been utilized to test performance under various privacy models. Common metrics include localization error, communication overhead, accuracy, runtime, and energy consumption. Some studies also account for noise effects [111] and time cost [116], reflecting a growing focus on the practical deployment and efficiency of these systems.
Many studies concentrate on defending against location privacy attacks and inference attacks [114], [118], [121]. Additionally, some works address probability-based and data privacy attacks [112], [123]. However, certain studies do not specify a particular attack model [109], [115], which may limit the general applicability of their privacy-preserving techniques.
There is a clear trend toward enhancing FL models by incorporating more sophisticated techniques. Earlier works utilized AutoEncoders with traditional privacy methods [107], while recent studies have integrated homomorphic encryption, federated knowledge distillation, and a mixture of experts to improve privacy without sacrificing performance [114], [121]. The growing use of deep learning models, such as CNNs, LSTMs, and GNNs, highlights a trend toward employing more complex architectures to achieve better performance and privacy preservation.
5.6.3 Discussion on employing existing FL-based techniques
To apply FL in ILF systems, the following recommendations can help maintain privacy while optimizing system performance:
Selecting the appropriate aggregation and privacy methods: When the primary concern is protecting against location privacy attacks, combining FedAvg with AutoEncoders offers an effective trade-off between security and computational efficiency. This approach ensures a reasonable level of privacy while maintaining system performance. However, for more complex threats, such as inference attacks or data privacy breaches, more sophisticated techniques may be necessary. Incorporating homomorphic encryption or federated knowledge distillation can significantly enhance the system’s resilience to such attacks, albeit at the cost of increased computational overhead. These advanced methods offer stronger privacy guarantees, particularly in scenarios where sensitive data is more vulnerable to sophisticated adversarial models.
Balancing performance and privacy: When performance metrics such as runtime and communication overhead are critical considerations, it is important to focus on methods that specifically address these concerns. Approaches like Monte Carlo Dropout and FedAvg are well-suited for maintaining a balance between high localization accuracy and reduced computational demands. By employing these techniques, it is possible to optimize system performance without significantly compromising on privacy or accuracy. These methods are particularly valuable in real-time applications, where minimizing time costs and communication overhead is essential for efficient system deployment and user experience.
Handling adversarial models: When a system is exposed to multiple types of attack models, including inference attacks or probability-based attacks, advanced privacy-preserving techniques become essential. Methods such as Federated Knowledge Distillation and homomorphic encryption are particularly well-suited to these scenarios, as they are designed to safeguard sensitive data against more sophisticated and adaptive adversaries. These techniques provide strong defenses while ensuring that system functionality and performance are not significantly compromised. By integrating these advanced methods, the system can maintain high levels of privacy and security, even in the face of evolving and complex threats, without sacrificing its overall utility or user experience.
5.7 Overall Comparison
In this section, first, we highlight key insights in ILF privacy preservation. Second, we explain a comparative assessment of ILFPPM techniques. Finally, we offer practical guidelines for ILFPPM deployment.
5.7.1 Key observations
The trade-off between privacy and performance in IPS highlights a significant challenge. Cryptographic techniques provide the strongest privacy guarantees but need high computational and communication overhead. In contrast, DP and FL offer a more practical balance between privacy and performance, though they may introduce localization errors or degrade QoS, reflecting the difficulty in optimizing both privacy and system efficiency.
Most privacy-preserving methods focus on location privacy attacks, while few address advanced attacks such as inference or data privacy attacks. Additionally, there is limited attention to fully malicious adversaries, especially within FL settings, indicating the need for more robust solutions to defend against sophisticated attacks.
The UJIIndoorLoc dataset is the most commonly used across all methods, serving as a standard benchmark for performance comparisons. However, this reliance on a single dataset limits the generalizability of solutions to diverse real-world scenarios. Cryptographic techniques, in particular, are often evaluated using private datasets or simulated environments, narrowing their applicability. Broader testing across varied datasets would improve the relevance and adaptability of these methods.
5.7.2 Comparative assessment of ILFPPM
Table 7: Comparative assessment of ILFPPM.
ILFPPM Device/Transmission/Server
Cryptography Transmission
Anonymization Device/Server
Differential Privacy Device/Server
Federated Learning Device